OpenStudy (anonymous):

Please help! There are some questions I need help solving, and some I want you to check..

OpenStudy (anonymous):

6.02] For the following system, if you isolated y in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 2x + y = 2 x − y = −6 −6 − x −6 + x 6 − x 6 + x

OpenStudy (anonymous):



OpenStudy (anonymous):

Or this... 6.03] Auditions for a new, upcoming band in Chicago were held with 72 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 34, how many women auditioned? 22 44 50 60

OpenStudy (anonymous):

6.05] There were 190 people at the Junior-Senior dance Sasha and her friends attended. Junior tickets were $2.00 each and Senior tickets were $3.50 each. The total receipts for the dance were $530. How many Juniors bought tickets? 70 80 90 100

OpenStudy (anonymous):

You can answer all 3, or just one, if u want. :)